# references.bib

### v1

@@ 1 Introduction
When the agent state satisfies the Markov property, i.e., $\mathds{P}(z_{t+1}\mid z_{1:t},a_{1:t})=\mathds{P}(z_{t+1}\mid z_{t},a_{t})$, where $a_{t}$ is the action, the optimal agent-state based policy can be obtained via a dynamic program (DP) [subramanian2022approximate]. An example of such an agent state is the belief state. However, the Markov property is not always satisfied, for example, it need not be satisfied when using frame stacking. In such settings, the best agent-state policy cannot be obtained via a DP. Nonetheless, there has been considerable interest to use RL to find a good agent-state based policy. One of the most commonly used RL algorithms is off-policy Q-learning, which we call agent-state Q-learning (ASQL). In ASQL for POMDPs, the Q-learning iteration is applied as if the agent state satisfied the Markov property even though it does not. The agent starts with an initial $Q_{1}(z,a)$, acts according to a behavior policy $\mu$, i.e., chooses $a_{t}\sim\mu(z_{t})$, and recursively updates
@@
where $R_{t}$ is the reward at time $t$ and the learning rates $\{\alpha_{t}\}_{t\geq 1}$ is chosen such that $\alpha_{t}(z,a)=0$ if $(z,a)\neq(z_{t},a_{t})$. The convergence of ASQL has been recently presented in [Kara2022, EWRL-RQL] which show that under some technical assumptions, ASQL converges to a limit. The policy determined by ASQL is the greedy policy w.r.t. this limit.
@@
As an illustration, consider the POMDP shown in Fig. 1, which is described in detail in App. A.2 as Ex. 2. Suppose the system starts in state $1$. Since the dynamics are deterministic, the agent can infer the current state from the history of past actions and can take the action to increment the current state and receive a per-step reward of $+1$. Thus, the performance $J^{\star}_{\textsc{{bd}}}$ of belief-state based policies is $J^{\star}_{\textsc{{bd}}}=1/(1-\gamma)$. Contrast this with the performance $J^{\star}_{\textsc{{sd}}}$ of deterministic agent-state base policies with agent state equal to current observation, which is given by $J^{\star}_{\textsc{{sd}}}=(1+\gamma-\gamma^{2})/(1-\gamma^{3})<J^{\star}_{\textsc{{bd}}}$.
@@
The above example highlights some salient features of periodic policies: (i) Periodic deterministic agent-state based policies may outperform stationary deterministic agent-state based policies. (ii) $\{J^{\star}_{L}\}_{L\geq 1}$is not a monotonically increasing sequence. This is because $\Pi_{L}$, the set of all periodic deterministic agent-state based policies of period $L$, is not monotonically increasing. (iii) If $L$ divides $M$, then $J^{\star}_{L}\leq J^{\star}_{M}$. This is because $\Pi_{L}\subseteq\Pi_{M}$. Therefore, periodic policies with period $L\in\{n!:n\in\mathds{N}\}$ will have monotonically increasing performance. (iv) In the above example, the set $\mathsf{D}_{0}$ is chosen such that the optimal sequence of actions [^1] Recall that the system dynamics are deterministic, so optimal policy can be implemented in open loop. is not periodic. Therefore, even though periodic policies can achieve a performance that is arbitrarily close to the optimal belief-based policies, they are not necessarily globally optimal (in the class of non-stationary agent-state based policies). Thus, the periodic deterministic policy class is a middle ground between the stationary deterministic and non-stationary policy classes and provides us a simple way of leveraging the benefits of non-stationarity while trading-off computational and memory complexity.
@@ 2.2 PASQL: Periodic agent-state based Q-learning algorithm for POMDPs
where the learning rate sequence $\{(\alpha^{0}_{t},\dots,\alpha^{L-1}_{t})\}_{t\geq 1}$ has the property that $\alpha^{\ell}_{t}(z,a)=0$ if $(\ell,z,a)\neq(\llbracket t\rrbracket,z_{t},a_{t})$. PASQL differs from ASQL in two aspects: (i) The behavior policy $\mu$ is periodic. (ii) The update of the Q-function is periodic. When $L=1$, PASQL collapses to ASQL.
@@ 2.3 Establishing the convergence of tabular PASQL
For all $(\ell,z,a)$, we have $\sum_{t\geq 1}\alpha^{\ell}_{t}(z,a)=\infty$ and $\sum_{t\geq 1}(\alpha^{\ell}_{t}(z,a))^{2}<\infty$.
@@
For any behavior policy $\mu$, the process $\{(S_{t},Y_{t},Z_{t},A_{t})\}_{t\geq 1}$ is a Markov chain.
@@
For the ease of notation, we will continue to use $\zeta_{\mu}^{\ell}$ to denote the marginal and conditional distributions w.r.t. $\zeta_{\mu}^{\ell}$. In particular, for marginals we use $\zeta_{\mu}^{\ell}(y,z,a)$ to denote $\sum_{s\in\mathsf{S}}\zeta_{\mu}^{\ell}(s,y,z,a)$ and so on; for conditionals, we use $\zeta_{\mu}^{\ell}(s|z,a)$ to denote $\zeta_{\mu}^{\ell}(s,z,a)/\zeta_{\mu}^{\ell}(z,a)$ and so on.
@@ 3 Numerical experiments
Convergence of PASQL with $L=2$. We assume that the agent state $Z_{t}=Y_{t}$ and take period $L=2$. We consider three behavioral policies: $\mu_{k}=(\mu_{k}^{0},\mu_{k}^{1})$, $k\in\mathsf{K}\coloneqq\{1,2,3\}$, where $\mu_{k}^{\ell}\colon\{0,1\}\to\Delta(\{0,1\})$, $\ell\in\{0,1\}$. The policy $\mu_{k}$ is completely characterized by four numbers which we write in matrix form as: $[\mu_{k}^{0}(0|0),\mu_{k}^{1}(0|1);\mu_{k}^{0}(0|1),\mu_{k}^{1}(0|1)]$. With this notation, the three policies are given by $\mu_{1}\coloneqq[0.2,0.8;0.2,0.8]$ , $\mu_{2}\coloneqq[0.5,0.5;0.5,0.5]$ , $\mu_{3}\coloneqq[0.8,0.2;0.8,0.2]$
@@
Advantage of learning periodic policies. As stated in the introduction, the main motivation of PASQL is that it allows us to learn non-stationary policies. To see why this is useful, we run ASQL (which is effectively PASQL with $L=1$). We again consider three behavioral policies: $\bar{\mu}_{k}$, $k\in\mathsf{K}\coloneqq\{1,2,3\}$, where $\bar{\mu}_{k}\colon\{0,1\}\to\Delta(\{0,1\})$, where (using similar notation as for $L=2$ case) $\bar{\mu}_{1}\coloneqq[0.2;0.2]$ , $\bar{\mu}_{2}\coloneqq[0.5;0.5]$ , $\bar{\mu}_{3}\coloneqq[0.8;0.8]$
@@ 5 Discussion
Non-stationary policies vs. memory augmentation. Non-stationarity is a fundamentally different concept than memory augmentation. As an illustration, consider the T-shaped grid world shown in Fig. 4, which has a corridor of length $2n$. In App. A.4, we show that for this example, a stationary policy which uses a sliding window of past $m$ observations and actions as the agent state needs a memory of at least $m>2n$ to reach the goal state. In contrast, a periodic policy with period $L=3$ can reach the goal state for every $n$. This example shows that periodicity is a different concept from memory augmentation and highlights the fact that mechanisms other than memory augmentation can achieve optimal behavior.
@@
Generalization to other variants. Our analysis is restricted to tabular off-policy Q-learning where a fixed behavioral policy is followed. Our proof fundamentally depends on the fact that the behavioral policy induces a cyclic limiting distribution on the periodic Markov chain $\{(S_{t},Y_{t},A_{t},Z_{t})\}_{t\geq 1}$. Such a condition is not satisfied in variants such as $\epsilon$-greedy Q-learning and SARSA. Generalizing the technical proof to cover these more practical algorithms (including function approximation) is an important future direction.
@@ B.1.1 Classification of states
    The states in a finite-state Markov chain can be partitioned into two sets: recurrent states and transient states. A state is recurrent if it is accessible from all states that is accessible from it (i.e., $i$ is recurrent if $i\rightsquigarrow j$ implies that $j\rightsquigarrow i$). States that are not recurrent are transient.
@@ Appendix D Stochastic Approximation with Markov noise
Consider the iterative process $\{X_{t}\}_{t\geq 1}$, where $X_{1}$ is arbitrary and for $t\geq 1$, we have
$X_{t+1}=(1-\alpha_{t})X_{t}+\alpha_{t}h(S_{t})$ (47)
@@
  • There is no “martingale noise” in (50), so A2.16 of [Bhatnagar2024] holds trivially.
@@
    Let $\tilde{h}_{\infty}(x,\nu)=\lim_{c\to\infty}h_{c}(x,\nu)==x$. Thus, the differential inclusion in A2.18(ii) is actually an ODE
@@
Fix any $\ell\in\mathsf{L}$, Consider the iterative process $\{X^{\ell}_{k}\}_{k\geq 1}$, where $X_{1}$ is arbitrary and for $k\geq 1$, we have
$X^{\ell}_{t+1}=(1-\alpha^{\ell}_{t})X^{\ell}_{t}+\alpha^{\ell}_{t}h(S_{t})$ (56)
@@
The high-level idea of the proof is similar to [Kara2022] for ASQL when the agent state is a finite window of past observations and action. The key observation of the [Kara2022] is the following: Consider an iterative process $X_{t+1}=(1-\alpha_{t})X_{t}+\alpha_{t}U_{t}$ with the learning rates $\alpha_{t}=1/(1+t)$. Then, $X_{t+1}=(X_{0}+\sum_{\tau=1}^{t}U_{t})/(1+t)$. Then, if the process $\{U_{t}\}_{t\geq 1}$ has an ergodic limit (e.g., when $\{U_{t}\}_{t\geq 1}$ is a function of a Markov chain, see Thm. 3), the process $\{X_{t}\}_{t\geq 1}$ converges to the ergodic limit of $\{U_{t}\}_{t\geq 1}$. We follow a similar idea but with the following changes:
@@ E.1 Step 1: State splitting of the error function
$\Delta^{\ell}_{t+1}(z,a)=(1-\alpha^{\ell}_{t}(z,a))\Delta^{\ell}_{t}(z,a)+\alpha^{\ell}_{t}(z,a)\bigl{[}U^{\ell,1}_{t}(z,a)+U^{\ell,2}_{t}(z,a)\bigr{]}$ (58)
@@
$U^{\ell,1}_{t}(z,a)$ $\coloneqq r(S_{t},A_{t})-r^{\ell}_{\mu}(z,a)+\gamma V_{\mu}^{\llbracket\ell+1\rrbracket}(Z_{t+1})-\gamma\sum_{z^{\prime}\in\mathsf{Z}}P^{\ell}_{\mu}(z^{\prime}|z,a)V_{\mu}^{\llbracket\ell+1\rrbracket}(z^{\prime}),$ (59)
@@
For each $\ell\in\mathsf{L}$, Eq. (58) may be viewed as a linear system with state $\Delta^{\ell}_{t+1}$ and two inputs $U^{\ell,1}_{t}$ and $U^{\ell,2}_{t}$. We exploit the linearity of the system and split the state into two components: $\Delta^{\ell}_{t+1}=X^{\ell,0}_{t+1}+X^{\ell,1}_{t+1}$, where the two components evolve as follows:
$X^{\ell,i}_{t+1}(z,a)=(1-\alpha^{\ell}_{t}(z,a))X^{\ell,i}_{t}(z,a)+\alpha^{\ell}_{t}(z,a)U^{\ell,i}_{t}(z,a),\quad i\in\{1,2\}$ (61)
@@
Linearity implies that (58) is equivalent to (61). We will now separately show that $\lVert X^{\ell,1}_{t}\rVert\to 0$ and $\lVert X^{\ell,2}_{t}\rVert\to 0$.
E.2 Step 2: Convergence of component $X^{\ell,1}_{t}$
Fix $(\ell,z,a)\in\mathsf{L}\times\mathsf{Z}\times\mathsf{A}$. Define the function

### v3

@@ 1 Introduction
When the agent state is an information state, i.e., satisfies the Markov property, i.e., $\mathds{P}(z_{t+1}|z_{1:t},a_{1:t})=\mathds{P}(z_{t+1}|z_{t},a_{t})$ and is sufficient for reward prediction, i.e., $\mathds{E}[R_{t}|y_{1:t},a_{1:t}]=\mathds{E}[R_{t}|z_{t},a_{t}]$ (where $y_{t}$ is the observation, $a_{t}$ is the action, and $R_{t}$ is the per-step reward), the optimal agent-state based policy can be obtained via a dynamic program (DP) [subramanian2022approximate]. An example of such an agent state is the belief state. But, in general, the agent state is not an information state. For example, frame stacking and RNN do not satisfy the Markov property, in general. It is also possible to have agent-states that satisfy the Markov property but are not sufficient for reward prediction (e.g., when the agent state is always a constant). In all such settings, the best agent-state policy cannot be obtained via a DP. Nonetheless, there has been considerable interest to use RL to find a good agent-state based policy.
One of the most commonly used RL algorithms is off-policy Q-learning, which we call agent-state Q-learning (ASQL). In ASQL for POMDPs, the Q-learning iteration is applied as if the agent state satisfied the Markov property even though it does not. The agent starts with an initial $Q_{1}(z,a)$, acts according to a behavior policy $\mu$, i.e., chooses $a_{t}\sim\mu(z_{t})$, and recursively updates
@@
where $\gamma\in[0,1)$ is the discount factor and the learning rates $\{\alpha_{t}\}_{t\geq 1}$ are chosen such that $\alpha_{t}(z,a)=0$ if $(z,a)\neq(z_{t},a_{t})$. The convergence of ASQL has been recently presented in [Kara2022, EWRL-RQL] which show that under some technical assumptions, ASQL converges to a limit. The policy determined by ASQL is the greedy policy w.r.t. this limit.
@@
As an illustration, consider the POMDP shown in Fig. 1, which is described in detail in App. A.2 as Ex. 2. Suppose the system starts in state $1$. Since the dynamics are deterministic, the agent can infer the current state from the history of past actions and can take the action to increment the current state and receive a per-step reward of $+1$. Thus, the performance $J^{\star}_{\textsc{{bd}}}$ of belief-state based policies is $J^{\star}_{\textsc{{bd}}}=1/(1-\gamma)$. Contrast this with the performance $J^{\star}_{\textsc{{sd}}}$ of deterministic agent-state base policies with agent state equal to current observation, which is given by $J^{\star}_{\textsc{{sd}}}=(1+\gamma-\gamma^{2})/(1-\gamma^{3})<J^{\star}_{\textsc{{bd}}}$. In particular, for $\gamma=0.9$, $J^{\star}_{\textsc{{bd}}}=10$ which is larger than $J^{\star}_{\textsc{{sd}}}=4.022$.
@@
The above example highlights some salient features of periodic policies: (i) Periodic deterministic agent-state based policies may outperform stationary deterministic agent-state based policies. (ii) $\{J^{\star}_{L}\}_{L\geq 1}$is not a monotonically increasing sequence. This is because $\Pi_{L}$, the set of all periodic deterministic agent-state based policies of period $L$, is not monotonically increasing. (iii) If $L$ divides $M$, then $J^{\star}_{L}\leq J^{\star}_{M}$. This is because $\Pi_{L}\subseteq\Pi_{M}$. In other words, if we take any integer sequence $\{L_{n}\}_{n\geq 1}$ that has the property that $L_{n}$ divides $L_{n+1}$, then the performance of the policies with period $L_{n}$ is monotonically increasing in $n$. For example, periodic policies with period $L\in\{n!:n\in\mathds{N}\}$ will have monotonically increasing performance. (iv) In the above example, the set $\mathsf{D}_{0}$ is chosen such that the optimal sequence of actions [^1] Recall that the system dynamics are deterministic, so optimal policy can be implemented in open loop. is not periodic. Therefore, even though periodic policies can achieve a performance that is arbitrarily close to the optimal belief-based policies, they are not necessarily globally optimal (in the class of non-stationary agent-state based policies). Thus, the periodic deterministic policy class is a middle ground between the stationary deterministic and non-stationary policy classes and provides us a simple way of leveraging the benefits of non-stationarity while trading-off computational and memory complexity.
@@ 2.2 PASQL: Periodic agent-state based Q-learning algorithm for POMDPs
where the learning rate sequence $\{(\alpha^{0}_{t},\dots,\alpha^{L-1}_{t})\}_{t\geq 1}$ is chosen such that $\alpha^{\ell}_{t}(z,a)=0$ if $(\ell,z,a)\neq(\llbracket t\rrbracket,z_{t},a_{t})$ and satisfies Assm. 1. PASQL differs from ASQL in two aspects: (i) The behavior policy $\mu$ is periodic. (ii) The update of the Q-function is periodic. When $L=1$, PASQL collapses to ASQL.
@@ 2.3 Establishing the convergence of tabular PASQL
For all $(\ell,z,a)$, the learning rates $\{\alpha^{\ell}_{t}(z,a)\}_{t\geq 1}$ are measurable with respect to the sigma-algebra generated by $(z_{1:t},a_{1:t})$ and satisfy $\alpha^{\ell}_{t}(z,a)=0$ if $(\ell,z,a)\neq(\llbracket t\rrbracket,z_{t},a_{t})$. Moreover, $\sum_{t\geq 1}\alpha^{\ell}_{t}(z,a)=\infty$ and $\sum_{t\geq 1}(\alpha^{\ell}_{t}(z,a))^{2}<\infty$, almost surely.
@@
For any behavior policy $\mu$, the process $\{(S_{t},Z_{t})\}_{t\geq 1}$ is Markov. Therefore, the processes $\{(S_{t},Z_{t},A_{t})\}_{t\geq 1}$ and $\{(S_{t},Y_{t},Z_{t},A_{t})\}_{t\geq 1}$ are also Markov.
@@
For the ease of notation, we will continue to use $\zeta_{\mu}^{\ell}$ to denote the marginal and conditional distributions w.r.t. $\zeta_{\mu}^{\ell}$. In particular, for marginals we use $\zeta_{\mu}^{\ell}(y,z,a)$ to denote $\sum_{s\in\mathsf{S}}\zeta_{\mu}^{\ell}(s,y,z,a)$ and so on; for conditionals, we use $\zeta_{\mu}^{\ell}(s|z,a)$ to denote $\zeta_{\mu}^{\ell}(s,z,a)/\zeta_{\mu}^{\ell}(z,a)$ and so on. Note that $\zeta_{\mu}^{\ell}(s,z,y,a)=\zeta_{\mu}^{\ell}(s,z)\mu(a\mid z)P(y|s,a)$. Thus, we have that $\zeta_{\mu}^{\ell}(s\mid z,a)=\zeta_{\mu}^{\ell}(s\mid z)$.
@@ 3 Numerical experiments
Convergence of PASQL with $L=2$. We assume that the agent state $Z_{t}=Y_{t}$ and take period $L=2$. We consider three behavioral policies: $\mu_{k}=(\mu_{k}^{0},\mu_{k}^{1})$, $k\in\mathsf{K}\coloneqq\{1,2,3\}$, where $\mu_{k}^{\ell}\colon\{0,1\}\to\Delta(\{0,1\})$, $\ell\in\{0,1\}$. The policy $\mu_{k}$ is completely characterized by four numbers which we write in matrix form as: $[\mu_{k}^{0}(0|0),\mu_{k}^{1}(0|0);\mu_{k}^{0}(0|1),\mu_{k}^{1}(0|1)]$. With this notation, the three policies are given by $\mu_{1}\coloneqq[0.2,0.8;0.8,0.2]$ , $\mu_{2}\coloneqq[0.5,0.5;0.5,0.5]$ , $\mu_{3}\coloneqq[0.8,0.2;0.2,0.8]$
@@
Advantage of learning periodic policies. As stated in the introduction, the main motivation of PASQL is that it allows us to learn non-stationary policies. To see why this is useful, we run ASQL (which is effectively PASQL with $L=1$). We again consider three behavioral policies: $\bar{\mu}_{k}$, $k\in\mathsf{K}\coloneqq\{1,2,3\}$, where $\bar{\mu}_{k}\colon\{0,1\}\to\Delta(\{0,1\})$, where (using similar notation as for $L=2$ case)
$\bar{\mu}_{1}\coloneqq[0.2;0.8]$ , $\bar{\mu}_{2}\coloneqq[0.5;0.5]$ , $\bar{\mu}_{3}\coloneqq[0.8;0.2]$
@@ 5 Discussion
Non-stationary policies vs. memory augmentation. Non-stationarity is a fundamentally different concept than memory augmentation. As an illustration, consider the T-shaped grid world (first considered in [NIPS2001_a38b1617]) shown in Fig. 4, which has a corridor of length $2n$. In App. A.4, we show that for this example, a stationary policy which uses a sliding window of past $m$ observations and actions as the agent state needs a memory of at least $m>2n$ to reach the goal state. In contrast, a periodic policy with period $L=3$ can reach the goal state for every $n$. This example shows that periodicity is a different concept from memory augmentation and highlights the fact that mechanisms other than memory augmentation can achieve optimal behavior.
@@
Generalization to other variants. Our analysis is restricted to tabular off-policy Q-learning where a fixed behavioral policy is followed. Our proof fundamentally depends on the fact that the behavioral policy induces a cyclic limiting distribution on the periodic Markov chain $\{(S_{t},Y_{t},Z_{t},A_{t})\}_{t\geq 1}$. Such a condition is not satisfied in variants such as $\epsilon$-greedy Q-learning and SARSA. Generalizing the technical proof to cover these more practical algorithms (including function approximation) is an important future direction.
Acknowledgments
The work of AS and AM was supported in part by a grant from Google’s Institutional Research Program in collaboration with Mila. The numerical experiments were enabled in part by support provided by Calcul Québec and Compute Canada.
@@ B.1.1 Classification of states
    The states in a finite-state Markov chain can be partitioned into two sets: recurrent states and transient states. A state is recurrent if it is accessible from all states that are from it (i.e., $i$ is recurrent if $i\rightsquigarrow j$ implies that $j\rightsquigarrow i$). States that are not recurrent are transient.
@@ Appendix D Stochastic Approximation with Markov noise
Let $\{M_{t+1}\}_{t\geq 1}$ be a square-integrable margingale difference sequence w.r.t. $\{\mathcal{F}_{t}\}_{t\geq 1}$ such that $\mathds{E}[M^{2}_{t+1}\mid\mathcal{F}_{t}]\leq K(1+\lVert X_{t}\rVert^{2})$ for some constant $K$. Consider the iterative process $\{X_{t}\}_{t\geq 1}$, where $X_{1}$ is arbitrary and for $t\geq 1$, we have
$X_{t+1}=(1-\alpha_{t})X_{t}+\alpha_{t}\bigl{[}h(S_{t})+M_{t+1}\bigr{]}.$ (47)
@@
  • The conditions on the martingale noise sequence $\{M_{t}\}_{t\geq 1}$ imply that A2.16 of [Bhatnagar2024] holds.
@@
    Let $\tilde{h}_{\infty}(x,\nu)=\lim_{c\to\infty}h_{c}(x,\nu)=x$. Thus, the differential inclusion in A2.18(ii) is actually an ODE
@@
Let $\{M_{t+1}\}_{t\geq 1}$ be a square-integrable margingale difference sequence w.r.t. $\{\mathcal{F}_{t}\}_{t\geq 1}$ such that $\mathds{E}[M^{2}_{t+1}\mid\mathcal{F}_{t}]\leq K(1+\lVert X_{t}\rVert^{2})$ for some constant $K$. Fix any $\ell\in\mathsf{L}$, Consider the iterative process $\{X^{\ell}_{k}\}_{k\geq 1}$, where $X_{1}$ is arbitrary and for $k\geq 1$, we have
$X^{\ell}_{t+1}=(1-\alpha^{\ell}_{t})X^{\ell}_{t}+\alpha^{\ell}_{t}\bigl{[}h(S_{t})+M_{t+1}\bigr{]}.$ (56)
@@
The high-level idea of the proof is similar to [Kara2022] for ASQL when the agent state is a finite window of past observations and action. The key observation of [Kara2022] is the following: Consider an iterative process $X_{t+1}=(1-\alpha_{t})X_{t}+\alpha_{t}U_{t}$ with the learning rates $\alpha_{t}=1/(1+t)$. Then, $X_{t+1}=(X_{0}+\sum_{\tau=1}^{t}U_{t})/(1+t)$. Then, if the process $\{U_{t}\}_{t\geq 1}$ has an ergodic limit (e.g., when $\{U_{t}\}_{t\geq 1}$ is a function of a Markov chain, see Thm. 3), the process $\{X_{t}\}_{t\geq 1}$ converges to the ergodic limit of $\{U_{t}\}_{t\geq 1}$. We follow a similar idea but with the following changes:
@@ E.1 Step 1: State splitting of the error function
$\Delta^{\ell}_{t+1}(z,a)=(1-\alpha^{\ell}_{t}(z,a))\Delta^{\ell}_{t}(z,a)+\alpha^{\ell}_{t}(z,a)\bigl{[}U^{\ell,0}_{t}(z,a)+U^{\ell,1}_{t}(z,a)+U^{\ell,2}_{t}(z,a)\bigr{]}$ (58)
@@
$U^{\ell,0}_{t}(z,a)$ $\coloneqq\left[r(S_{t},A_{t})-r^{\ell}_{\mu}(z,a)\right]\mathds{1}_{\{Z_{t}=z,A_{t}=a\}},$ (59)
$U^{\ell,1}_{t}(z,a)$ $\coloneqq\Bigl{[}\gamma V_{\mu}^{\llbracket\ell+1\rrbracket}(Z_{t+1})-\gamma\sum_{z^{\prime}\in\mathsf{Z}}P^{\ell}_{\mu}(z^{\prime}|z,a)V_{\mu}^{\llbracket\ell+1\rrbracket}(z^{\prime})\Bigr{]}\mathds{1}_{\{Z_{t}=z,A_{t}=a\}},$ (60)
@@
Note that we have added extra indicator functions in the $U^{\ell,i}_{t}(z,a)$ terms, $i\in\{0,1\}$. This does not change the value of $\alpha^{\ell}_{t}(z,a)U^{\ell,i}_{t}(z,a)$ because the learning rates have the property that $\alpha^{\ell}_{t}(z,a)=0$ if $(\ell,z,a)\neq(\llbracket t\rrbracket,z_{t},a_{t})$ (see Assm. 1).
@@
For each $\ell\in\mathsf{L}$, Eq. (58) may be viewed as a linear system with state $\Delta^{\ell}_{t+1}$ and three inputs $U^{\ell,0}_{t}$,$U^{\ell,1}_{t}$ and $U^{\ell,2}_{t}$. We exploit the linearity of the system and split the state into three components: $\Delta^{\ell}_{t+1}=X^{\ell,0}_{t+1}+X^{\ell,1}_{t+1}+X^{\ell,2}_{t+1}$, where the three components evolve as follows:
$X^{\ell,i}_{t+1}(z,a)=(1-\alpha^{\ell}_{t}(z,a))X^{\ell,i}_{t}(z,a)+\alpha^{\ell}_{t}(z,a)U^{\ell,i}_{t}(z,a),\quad i\in\{0,1,2\}$ (62)
@@
Linearity implies that (58) is equivalent to (62). We will now separately show that $\lVert X^{\ell,0}_{t}\rVert\to 0$, $\lVert X^{\ell,1}_{t}\rVert\to 0$ and $\lVert X^{\ell,2}_{t}\rVert\to 0$.
E.2 Step 2: Convergence of component $X^{\ell,0}_{t}$
Fix $(\ell,z_{\circ},a_{\circ})\in\mathsf{L}\times\mathsf{Z}\times\mathsf{A}$ and define
$h_{r}(S_{t},Z_{t},A_{t};\ell,z_{\circ},a_{\circ})$ $=\bigl{[}r(S_{t},A_{t})-r^{\ell}_{\mu}(z_{\circ},a_{\circ})\bigr{]}\mathds{1}_{\{Z_{t}=z_{\circ},A_{t}=a_{\circ}\}}.$
Then the process $\{X^{\ell,0}_{t}(z_{\circ},a_{\circ})\}_{t\geq 1}$ is given by the stochastic iteration
$X^{\ell,0}_{t+1}(z_{\circ},a_{\circ})=(1-\alpha^{\ell}_{t}(z_{\circ},a_{\circ}))X^{\ell,0}_{t}(z_{\circ},a_{\circ})+\alpha^{\ell}_{t}(z_{\circ},a_{\circ})h_{r}(S_{t},Z_{t},A_{t};\ell,z_{\circ},a_{\circ}),$ (63)
which is of the form (56). The process $\{(S_{t},Z_{t},A_{t})\}_{t\geq 1}$ is a periodic Markov chain and the learning rates $\{\alpha^{\ell}_{t}(z_{\circ},a_{\circ})\}_{t\geq 1}$ satisfy the conditions of Prop. 4 due to Assm. 1. Therefore, Prop. 4 implies that $\{X^{\ell,0}_{t}(z_{\circ},a_{\circ})\}_{t\geq 1}$ converges a.s. to the following limit
$\lim_{t\to\infty}X^{\ell,0}_{t}(z_{\circ},a_{\circ})$ $=\sum_{s,z,a\in\mathsf{S}\times\mathsf{Z}\times\mathsf{A}}\zeta_{\mu}^{\ell}(s,z,a)h_{r}(s,z,a;\ell,z_{\circ},a_{\circ})$
$=\sum_{s,z,a\in\mathsf{S}\times\mathsf{Z}\times\mathsf{A}}\zeta_{\mu}^{\ell}(s,z,a)\mathds{1}_{\{z=z_{\circ},a=a_{\circ}\}}\bigl{[}r(s,a)-r^{\ell}_{\mu}(z_{\circ},a_{\circ})\bigr{]}$
$=\biggl{[}\sum_{s\in\mathsf{S}}\zeta_{\mu}^{\ell}(s,z_{\circ},a_{\circ})r(s,a_{\circ})\biggr{]}-\zeta_{\mu}^{\ell}(z_{\circ},a_{\circ})r^{\ell}_{\mu}(z_{\circ},a_{\circ})$
$=\biggl{[}\sum_{s\in\mathsf{S}}\zeta_{\mu}^{\ell}(s,z_{\circ},a_{\circ})r(s,a_{\circ})\biggr{]}-\biggl{[}\sum_{s\in\mathsf{S}}\zeta_{\mu}^{\ell}(z_{\circ},a_{\circ})\zeta_{\mu}^{\ell}(s|z_{\circ})r(s,a_{\circ})\biggr{]}$
$=\biggl{[}\sum_{s\in\mathsf{S}}\zeta_{\mu}^{\ell}(s,z_{\circ})\mu(a_{\circ}|z_{\circ})r(s,a_{\circ})\biggr{]}-\biggl{[}\sum_{s\in\mathsf{S}}\zeta_{\mu}^{\ell}(z_{\circ})\mu(a_{\circ}|z_{\circ})\zeta_{\mu}^{\ell}(s|z_{\circ})r(s,a_{\circ})\biggr{]}$
$=0$
Hence, for all $(\ell,z_{\circ},a_{\circ})$, the process $\{X^{\ell,0}_{t}(z_{\circ},a_{\circ})\}_{t\geq 1}$ converges to zero almost surely.
E.3 Step 3: Convergence of component $X^{\ell,1}_{t}$
Let $W_{t}$ denote the tuple $(S_{t},Z_{t},A_{t},S_{t+1},Z_{t+1},A_{t+1})$. Note that $\{W_{t}\}_{t\geq 1}$ is also a periodic Markov chain and converges to a cyclic limiting distribution $\bar{\zeta}^{\ell}_{\mu}$, where
$\bar{\zeta}^{\ell}_{\mu}(s,z,a,s^{\prime},z^{\prime},a^{\prime})=\zeta_{\mu}^{\ell}(s,z,a)\sum_{y^{\prime}\in\mathsf{Y}}P(s^{\prime},y^{\prime}|s,a)\mathds{1}_{\{z^{\prime}=\phi(z,y^{\prime},a)\}}\mu(a^{\prime}|z^{\prime}).$
We use $\bar{\zeta}_{\mu}^{\ell}(s,z,a,\mathsf{S},\mathcal{Z},\mathcal{A})$ to denote the marginalization over the “future states” and a similar notation for other marginalizations. Note that $\bar{\zeta}_{\mu}^{\ell}(s,z,a,\mathsf{S},\mathcal{Z},\mathcal{A})=\zeta_{\mu}^{\ell}(s,z,a)$.
Fix $(\ell,z_{\circ},a_{\circ})\in\mathsf{L}\times\mathsf{Z}\times\mathsf{A}$ and define